A rainstorm in Portland, Oregon, wiped out the electricity in 7% of the households in the city. Suppose that a random sample of 50 Portland households is taken after the rainstorm.
A Estimate the number of households in the sample that lost electricity by giving the mean of the relevant distribution (that is, the expectation of the relevant random variable). Do not round your response.
B Quantify the uncertainty of your estimate by giving the standard deviation of the distribution. Round your response to at least three decimal places.

Answer by stanbon(75887)

A rainstorm in Portland, Oregon, wiped out the electricity in 7% of the households in the city. Suppose that a random sample of 50 Portland households is taken after the rainstorm.
-------------------------------------------------
A Estimate the number of households in the sample that lost electricity by giving the mean of the relevant distribution (that is, the expectation of the relevant random variable). Do not round your response.
---
The distribution is binomial (households either do or do not lose electricity).
Mean = np = 50*0.07 = 3.5 households
--------------------------------------------------------------
B Quantify the uncertainty of your estimate by giving the standard deviation of the distribution. Round your response to at least three decimal places
---
s = sqrt(npq) = sqrt(50*0.07*0.93) = 1.804..
-------------------------------------------------
Cheers,
Stan H.